OpenStudy (anonymous):

how do i convert r^2 sin(2 theta) to cartesian form ?

OpenStudy (amoodarya):

x=r cos theta y=r sin theta x^2+y^2=r^2 so r^2 sin(2 theta) = r^2 2 sin(theta) cos (theta) =2 r cos (theta) r sin (theta) =2xy

OpenStudy (amoodarya):

welcome to open study

OpenStudy (amoodarya):